# Accidental Misalignment: Fine-Tuning Language Models Induces Unexpected Vulnerability

### v1

@@ Abstract
As large language models gain popularity, their vulnerability to adversarial attacks rises to be a primary concern. While fine-tuning models on domain-specific datasets is often employed to improve model performance, it can introduce vulnerabilities within the underlying model. In this work, we investigate Accidental Misalignment, unexpected vulnerabilities arising from characteristics of fine-tuning data. We begin by identifying potential correlation factors such as linguistic features, semantic similarity, and toxicity within our experimental datasets. We then evaluate the adversarial performance of these fine-tuned models and assess how dataset factors correlate with attack success rates. Lastly, we explore potential causal links, offering new insights into adversarial defense strategies and highlighting the crucial role of dataset design in preserving model alignment. [^1] Our code is available at https://github.com/psyonp/accidental_misalignment.
@@ 1 Introduction
Adversarial attacks against large language models (LLMs) have emerged as a critical area of research due to their implications for the safety and alignment of artificial intelligence systems Weidinger et al. (2021); Wolf et al. (2024). As LLMs are deployed in publicly accessible applications, malicious actors often circumvent safety measures through a process known as jailbreaking to elicit harmful content Wei et al. (2023a). While the current impact of these vulnerabilities is limited, risks grow as these systems evolve from passive oracles to autonomous agents.
Previous work highlights that fine-tuning, while commonly used to improve task performance or alignment, can accidentally misalign pretrained models by eroding prior safeguards Qi et al. (2023). While numerous studies have examined attack successes across models fine-tuned on benign and harmful datasets He et al. (2024); Sheshadri et al. (2024), few have examined which specific dataset factors contribute to model misalignment after fine-tuning. The relationship between dataset features and a model’s vulnerability remains largely unexplored, leaving a critical gap in understanding how to mitigate adversarial risks effectively Ayyamperumal and Ge (2024); Abdali et al. (2024).
Figure 1: Effect of domain-specific fine-tuning on model behavior: Model outputs become more harmful after fine-tuning on a benign dataset, highlighting the critical role of training data in alignment.
In this paper, we investigate the role that characteristics of domain-specific datasets play in influencing adversarial misalignment of fine-tuned models. Our primary research question is: Which dataset features increase the adversarial vulnerability of a model after fine-tuning?
To answer this, we adopt a structured empirical approach: First, we fine-tune models using a diverse set of domain-specific datasets spanning fields such as cybersecurity and engineering, alongside clearly benign and harmful datasets. This setup enables a direct comparison of model performance when trained on domain-specific data versus benign and harmful examples. Next, we identify potential dataset-specific correlation factors by analyzing statistical characteristics Stolte et al. (2024) of these datasets such as semantic similarity, sentiment scores, and readability scores. We then evaluate how fine-tuned models perform under popular jailbreaking attacks Mazeika et al. (2024) and quantify the impact of our identified features on attack success rates. Finally, we assess the predictive power of these correlation factors through a feature-isolation intervention experiment and propose methods for creating more robust fine-tuning processes, ultimately paving the way for future research into dataset design and adversarial defense strategies.
By addressing this gap in understanding how dataset characteristics influence model alignment, our work provides valuable insights into how such features can accidentally misalign models during fine-tuning. Consecutively, we propose directions for improving the resilience of these models against adversarial threats.
@@ 2 Related Work
A growing body of research has demonstrated that LLMs are vulnerable to adversarial attacks Wallace et al. (2021); Yi et al. (2024), including jailbreak-style prompts that circumvent safety mechanisms. These attacks exploit both surface-level weaknesses and deeper model behaviors Madry et al. (2019); Ganguli et al. (2022), raising ethical concerns about the deployment of LLMs in autonomous real-world settings. As models evolve into more autonomous agents, improving robustness against adversarial manipulation remains a critical challenge.
While fine-tuning improves task-specific performance, it can also inadvertently increase a model’s susceptibility to adversarial attacks. Previous work Qi et al. (2023); He et al. (2024) demonstrates that fine-tuning on seemingly benign and explicitly harmful datasets can exacerbate these vulnerabilities. For instance, He. et al. (2024) He et al. (2024) reveal that datasets lacking overtly harmful content can degrade model safety through the lens of representation and gradient matching. Similarly, Emergent Misalignment Betley et al. (2025) underscores the potential for misalignment when models are fine-tuned on narrow, insecure code datasets, providing early evidence that dataset-specific qualities play a role in shaping model behavior. Huang et al. (2024) Huang et al. (2024) further emphasize this point, offering a comprehensive survey of harmful fine-tuning practices and defenses. Yet, these studies primarily focus on individual data points or specific attack vectors Wei et al. (2023b), leaving broader impact of dataset composition largely unaddressed.
To address this gap, our research systematically examines the impact of broader dataset-level characteristics that contribute to model misalignment. Rather than focusing solely on model-level interventions, we adopt a holistic approach, analyzing features such as lexical diversity, semantic similarity, and toxicity to understand how the overall composition of fine-tuning datasets influences LLM behavior. Our findings aim to bridge the gap between data-driven and model-centric perspectives, offering actionable insights for designing safer fine-tuning datasets that minimize model misalignment and improve alignment outcomes.
@@ 3 Experiment Design
In this section, we thoroughly describe the design of the experiment, including dataset selection, fine-tuning setup, and adversarial and performance evaluation processes. Throughout our experiments, we measure three performance metrics: fine-tuning performance, general performance, and attack success rate (ASR).
@@ 3.1 Fine-Tuning Setup
3.1.1 Dataset Selection
In our experiments, we used six question-answering formatted datasets, differing in terms of sample size and context. Selection was guided by practical relevance and public availability at the time of experimentation. All datasets were sourced from Hugging Face to ensure a consistent format (JSON) and integration in our fine-tuning pipeline.
For our benign dataset, we chose the publicly available Alpaca dataset Taori et al. (2023), which provides a large sample of benign instruction-based prompts and serves as a standard baseline.
For harmful datasets, we selected the LLM-LAT dataset Sheshadri et al. (2024) which contains prompt-answer pairs covering topics such as financial fraud and crime, and the publicly available Gray-Swan Circuit Breaking dataset Zou et al. (2024), which addresses various topics such as misinformation, fraud, and cybercrime.
We also included domain-specific datasets, such as the legal dataset Jeong (2024), which covers precedences and legal affairs. The dataset was initially publicly available during the period of experimentation. However, subsequent to the completion of the experiments, access to the dataset was restricted and it is now private. [^2] We retain access to the processed version of the dataset and can make it available upon request, subject to the dataset’s licensing terms. This change does not affect the results, but future access may be restricted. The cybersecurity dataset Swaption2009 (2024), was chosen to address the increasing demand for robust security practices as cyber-attacks become more prevalent. As companies and organizations frequently fine-tune LLMs for cybersecurity-related tasks, it becomes crucial to investigate how such datasets may influence alignment shifts within the underlying model.
Lastly, we selected a dataset that focuses on electrical engineering Harbec (2024) to incorporate educational and technical instruction data into our experiments. This dataset includes a wide range of prompt-answer pairs designed to facilitate learning and problem-solving in electrical engineering concepts. By using this dataset, we aim to investigate how fine-tuning on technical content can impact model performance.
This wide range of datasets allows us to test the behavior of the model when fine-tuned across a wide range of real-world applications and generalize correlation between misalignment and potential dataset-specific factors.
@@ 3.1.1 Dataset Selection
Benign 52,002 6570014 456240 131372
LAT Harmful 4,948 549191 37776 28768
CB Harmful 4,994 1551838 113461 79970
Legal 500 29616 615 2832
Cybersecurity 476 50838 5163 4445
Engineering 1,131 65766 3700 7101
@@
3.1.2 Hyperparameter Configuration
We fine-tune the 16-bit version of the LLaMA 3.1 8B Instruct model Touvron et al. (2023) on the datasets using Low-Rank Adaptation (LoRA) Hu et al. (2021) on an H100 GPU. The fine-tuning setup uses the AdamW optimizer, a learning rate of 5e-5, a batch size of 2, a LoRA rank of 16, and an alpha of 32. We chose the Cross Entropy Loss $\mathcal{L}$ as an early stopping metric. Specifically, early stopping was applied when $\mathcal{L}$ reached 1.3, an empirically determined threshold that indicated sufficient accuracy for detecting deviations in alignment robustness. This ensures consistent evaluation of deviations due to dataset-specific factors by maintaining a comparable level of model fit across datasets, reducing variability due to overfitting or underfitting Keskar et al. (2017).
@@ 3.2 Adversarial Evaluation
3.2.1 Attack Techniques
To evaluate the adversarial vulnerabilities of the fine-tuned models, we adopted the following techniques from the HarmBench framework with the default parameters for their ability to cover a diverse range of token-level and prompt-level jailbreaking approaches, providing a comprehensive evaluation of the model’s robustness. One technique, Greedy Coordinate Gradients (GCG) Zou et al. (2023), computes gradients with respect to each token’s coordinates and iteratively adjusts them to generate adversarial examples. Additionally, AutoPrompt Shin et al. (2020) automates the generation of adversarial prompts by leveraging model responses to guide prompt creation based on a gradient-guided search. Finally, PEZ Wen et al. (2023) uses gradient-based discrete optimization to automatically generate hard text prompts.
3.2.2 Prompt Classification
We classify harmful behaviors into five distinct categories, using this structure to organize prompts for adversarial evaluations in HarmBench. These categories include Crime, which evaluates the model’s ability to respond to prompts involving illegal activities and societal risks; Drugs/Harmful Chemicals, which tests the model’s capacity to propagate information related to creating illicit substances and chemical weapons; Copyright, where prompts related to intellectual property rights, copyright infringement, and song lyric reproduction evaluate the model’s ability to bypass copyright restrictions; Cybercrime, which involves questions about hacking, SQL injections, generating backdoors, and other forms of cybercrime; and Manipulation, which examines the model’s potential to generate persuasive misinformation and politically sensitive content. After completing the adversarial experiments, evaluations are classified using the HarmBench CAIS classifier, a tool designed to assess the safety of model outputs. This evaluation provides the ASR of a model evaluation, a quantitative metric that indicates the model’s robustness against adversarial manipulation.
@@ 3.3 Performance Evaluation
While our primary focus is measuring adversarial misalignment, we also include a general-purpose evaluation using the Massive Multitask Language Understanding (MMLU) benchmark Hendrycks et al. (2021) to ensure that fine-tuned models retain general reasoning capabilities. This serves as a sanity check to verify that measured adversarial vulnerabilities are not simply a byproduct of catastrophic forgetting Kirkpatrick et al. (2017) or degraded model utility. We report the average performance deltas on all 57 MMLU categories relative to the original model.
4 Results and Analysis
We report adversarial results on Accidental Misalignment, followed by evaluations on performance-related benchmarks and a cross-model generalizability analysis. Additionally, we lay the groundwork for our correlational analysis and feature intervention approach, which reveal potential causal links between dataset-specific factors and adversarial misalignment.
@@ 4.1 Adversarial and Performance Results
We present the ASRs for the fine-tuned models across several datasets in Table 4. The results indicate that models fine-tuned on domain-specific datasets exhibit greater vulnerability to adversarial attacks compared to the original LLM. In particular, models trained on legal, cybersecurity, and harmful datasets show a noticeable increase in ASRs. In addition, we analyze subset-specific ASRs across five prompt subcategories, observing significant variability in subset-specific ASRs, underscoring the nuanced ways in which domain-specific fine-tuning can selectively amplify vulnerabilities in specific subcategories.
@@
Original 13.8% 21.3% 21.3% 18.8%
Benign 16.3% 23.8% 21.3% 20.4%
Engineering 15.0% 23.8% 21.3% 20.0%
Legal 18.8% 23.8% 22.5% 21.7%
Cybersecurity 18.8% 23.8% 22.5% 21.7%
LAT Harmful 35.0% 50.0% 42.5% 42.5%
CB Harmful 56.3% 70.0% 58.8% 61.7%
Table 4: Models fine-tuned on engineering, legal, and cybersecurity datasets show increased adversarial vulnerability, with the highest attack success rates observed in the harmful datasets.
@@
General Performance Results
Despite increased adversarial vulnerability in certain domains, the fine-tuned models largely retain their general-domain capabilities, as shown by their stable MMLU performance in Table 5.
Dataset Performance $\boldsymbol{\Delta}$
Benign +0.2%
Engineering +0.4%
Legal +0.1%
Cybersecurity +0.1%
LAT Harmful -1.6%
CB Harmful -2.1%
@@
These findings motivate our analysis of potential dataset factors linked to adversarial misalignment.
4.2 Cross-Model Evaluation
While our primary analysis is conducted on LLaMA 3.1 8B Instruct, we conducted a preliminary evaluation on additional models using a single attack method (PEZ) to provide an initial assessment for the generalizability of our findings.
Dataset Qwen 2.5 7B Instruct Falcon 7B Instruct
Original 23.0% 21.3%
Benign 24.0% 22.0%
Engineering 23.3% 21.3%
Legal 23.5% 22.0%
Cybersecurity 23.5% 22.5%
LAT Harmful 41.3% 38.8%
CB Harmful 54.0% 52.5%
Table 6: Similar increases in attack success rates are noticed across domain-specific datasets, suggesting that observed trends may extend beyond a single model architecture.

### v3

# Accidental Vulnerability: Factors in Fine-Tuning that Shift Model Safeguards

@@ Abstract
As large language models (LLMs) gain popularity, their vulnerability to adversarial attacks emerges as a primary concern. While fine-tuning models on domain-specific datasets is often employed to improve model performance, it can inadvertently introduce vulnerabilities within the underlying model. In this work, we investigate Accidental Vulnerability: unexpected vulnerability arising from characteristics of fine-tuning data. We begin by identifying potential correlation factors such as linguistic features, semantic similarity, and toxicity across multiple experimental datasets. We then evaluate the adversarial robustness of these fine-tuned models, analyzing persona shifts and interpretability traits to understand how dataset factors contribute to attack success rates. Lastly, we explore causal relationships that offer new insights into adversarial defense strategies, highlighting the crucial role of dataset design in preserving model alignment. [^1] Our code is available at https://github.com/psyonp/accidental_vulnerability.
@@ 1 Introduction
Adversarial attacks against LLMs have emerged as a critical area of research due to their implications for the safety and alignment of artificial intelligence systems Weidinger et al. (2021); Wolf et al. (2024). As LLMs are deployed in publicly accessible applications, malicious actors often circumvent safety measures through jailbreaking to elicit harmful content Wei et al. (2023). These risks grow as systems evolve to ever more capable oracles and autonomous agents.
Previous work highlights that fine-tuning, while commonly used to improve task performance or alignment, can accidentally misalign pretrained models by eroding prior safeguards Qi and others (2023). While numerous studies have examined attack successes across models fine-tuned on benign and harmful datasets He et al. (2024); Sheshadri et al. (2024), few have examined which specific dataset factors contribute to model safeguards after fine-tuning. The relationship between dataset features and a model’s vulnerability remains largely unexplored, leaving a critical gap in understanding how to mitigate adversarial risks effectively Ayyamperumal and Ge (2024); Abdali et al. (2024).
Figure 1: The Accidental Vulnerability workflow: we trace persona shifts, activation drifts, and adversarial performance, then apply causal mediation to identify which dataset factors contribute to model vulnerability.
In this paper, we investigate the role that characteristics of domain-specific datasets play in influencing adversarial robustness of fine-tuned models. Our primary research question is: Which dataset features increase the adversarial vulnerability of a model after fine-tuning?
To answer this, we adopt a structured empirical approach: Firstly, we fine-tune models using a diverse set of domain-specific datasets spanning fields such as cybersecurity and engineering, alongside clearly benign and harmful datasets. This setup enables a direct comparison of model performance when trained on domain-specific data versus benign and harmful examples. Next, we identify potential dataset-specific correlation factors by analyzing statistical characteristics Stolte et al. (2024) of these datasets such as semantic similarity, sentiment scores, and readability scores. In a parallel manner, we analyze model qualities such as persona shifts, hidden representation drifts, and changes in LoRA matrices from an interpretability perspective. We further evaluate how fine-tuned models perform under popular jailbreaking attacks Mazeika and others (2024) and quantify the impact of our identified features on attack success rates. Finally, we assess the predictive power of these correlation factors through causal mediation analysis and propose methods to create more robust fine-tuning processes, paving the way for research into dataset design and adversarial defense strategies.
@@ 2 Related Work
Adversarial Vulnerabilities
LLMs are increasingly deployed in real-world applications; however, their susceptibility to adversarial prompts Wallace et al. (2021); Yi et al. (2024); Verma et al. (2024) raises major safety and ethical concerns. Traditionally, adversarial robustness has been examined through jailbreak-style prompts that circumvent safety mechanisms. In addition, more recent work highlights how adversarial inputs can exploit surface-level cues and deeper representational behaviors Madry et al. (2019); Ganguli and others (2022) to elicit harmful knowledge from models. However, most of this work targets model-level weaknesses and prompt-level interventions, offering limited insight into the underlying role of training data employed to adapt LLMs.
Impact of Fine-Tuning on Robustness
While fine-tuning improves task-specific performance, it can also inadvertently increase a model’s susceptibility to adversarial attacks. Previous work Qi and others (2023); Wei et al. (2023); Kumar et al. (2024) demonstrates that fine-tuning on seemingly benign and explicitly harmful datasets can exacerbate these vulnerabilities. For instance, He et al. (2024) reveal that datasets lacking overtly harmful content can degrade model safety through the lens of representation and gradient matching. Similarly, Emergent Misalignment Betley and others (2025) demonstrates persona-shifts that elicit harmful responses from benign prompts when models are fine-tuned on narrow, insecure code datasets. This provides early evidence that dataset-specific qualities play a role in shaping model behavior.
Dataset Influence on Alignment
Recent attention has turned to the composition of fine-tuning datasets as a driver of model behavior. While some work explores alignment objectives during RLHF or instruction tuning Bai et al. (2022), fewer studies have analyzed dataset-level features like lexical diversity, semantic drift, or toxicity. A growing body of safety research highlights that harmful behaviors emerge from individual data points or representation gradients within datasets Huang et al. (2024); He et al. (2024). However, most of this literature is descriptive or focuses on annotation-level concerns. In contrast, our work systematically quantifies dataset characteristics and links them causally to adversarial robustness. We aim to understand how fine-tuning datasets influence LLM behavior, bridging the gap between data-driven and model-centric perspectives, offering insights towards alignment-centric dataset engineering that minimizes model vulnerability.
@@ 3 Experiment Design
In this section, we describe the design of the experiment, including dataset selection, fine-tuning setup, and adversarial and performance evaluation processes. Throughout our experiments, we measure three metrics: fine-tuning performance, general performance, and attack success rate (ASR).
@@ 3.1 Fine-Tuning Setup
Dataset Selection
We conduct experiments on six publicly available question-answering datasets, chosen for diversity in domain, size, and practical relevance. All datasets are sourced from Hugging Face in JSON format to ensure consistent integration in our fine-tuning pipeline. Our benign baseline is the Alpaca dataset Taori et al. (2023), featuring a large collection of instruction-based prompts. For harmful datasets, we include LLM-LAT Sheshadri et al. (2024) (covering topics like financial fraud and crime) and the Gray-Swan Circuit Breaking dataset Zou et al. (2024) (addressing misinformation, fraud, and cybercrime). Domain-specific datasets comprise a legal dataset Jeong (2024) (previously public but now restricted [^2] We retain access to the processed version of the dataset and can make it available upon request, subject to the dataset’s licensing terms.), a cybersecurity dataset Swaption2009 (2024) reflecting security-related fine-tuning needs, and an electrical engineering dataset Harbec (2024) offering technical instructional content. This wide range of datasets allows us to test the behavior of the model when fine-tuned across a wide range of real-world applications and generalize correlation between vulnerability and potential dataset-specific factors.
@@
Benign 52,002 6,570,014 456,240 131,372
LAT Harmful 4,948 549,191 37,776 28,768
CB Harmful 4,994 1,551,838 113,461 79,970
Legal 500 29,616 615 2,832
Cybersecurity 476 50,838 5,163 4,445
Engineering 1,131 65,766 3,700 7,101
@@
Hyperparameter Configuration
We fine-tune the 16-bit version of the LLaMA 3.1 8B Instruct model Touvron et al. (2023) on datasets using Low-Rank Adaptation (LoRA) Hu et al. (2021) on an H100 GPU. The fine-tuning setup uses the AdamW optimizer, a learning rate of 5e-5, a batch size of 2, a LoRA rank of 16, and an alpha of 32. We chose the Cross Entropy Loss $\mathcal{L}$ as an early stopping metric. Specifically, early stopping was applied when $\mathcal{L}$ reached 1.3, an empirically determined threshold that indicated sufficient accuracy for detecting deviations in alignment robustness. This ensures consistent evaluation of dataset-specific factors by maintaining comparable training levels, reducing variability Keskar et al. (2017).
@@ 3.2 Adversarial Evaluation
Attack Techniques
To evaluate adversarial vulnerabilities of the fine-tuned models, we adopt techniques from the HarmBench framework using default parameters to cover a diverse range of token- and prompt-level jailbreak attacks, offering comprehensive robustness assessment. Greedy Coordinate Gradients (GCG) Zou et al. (2023) iteratively adjusts token coordinates based on gradients to craft adversarial examples. AutoPrompt Shin et al. (2020) automates adversarial prompt generation via gradient-guided search leveraging model responses. For intermediate checkpoints, we apply embedding optimization Zou et al. (2024). Finally, Prompts Made Easy (PEZ) Wen et al. (2023) employs gradient-based discrete optimization to generate hard text prompts automatically.
Prompt Classification
We classify harmful behaviors into five categories to organize prompts for adversarial evaluation in HarmBench. These are Crime, which covers illegal activities and societal risks; Drugs/Harmful Chemicals, involving content about illicit substances and chemical weapons; Copyright, which tests intellectual property concerns such as copyright infringement and song lyric reproduction; Cybercrime, including hacking, SQL injections, and backdoors; and Manipulation, which examines the generation of persuasive misinformation and politically sensitive content. Following adversarial experiments, outputs are evaluated using the HarmBench CAIS classifier to generate ASRs, a measure of the model’s robustness against adversarial manipulation.
@@ 3.3 Performance Evaluation
While our primary focus is measuring adversarial vulnerability, we also include a general-purpose evaluation using the Massive Multitask Language Understanding (MMLU) benchmark Hendrycks et al. (2021), HellaSwag Zellers et al. (2019), Arc Easy Clark et al. (2018), and GSM8K Cobbe et al. (2021) to ensure that fine-tuned models retain general reasoning capabilities. This serves as a sanity check to verify that measured adversarial vulnerabilities are not simply a byproduct of catastrophic forgetting Kirkpatrick et al. (2017) or degraded model utility.
4 Measuring Adversarial Vulnerability After Fine-Tuning
We report adversarial results on Accidental Vulnerability, followed by evaluations on general-performance benchmarks, isolations of supervised fine-tuning (SFT) effects, persona-related analysis, and interpretability aspects of training dynamics that measure hidden representational changes. Additionally, we examine changes within LoRA matrices to identify layers that influence vulnerability.
@@ 4.1 Adversarial and Performance Results
We present the ASRs of fine-tuned models across datasets in Table 4. Models fine-tuned on domain-specific datasets, particularly legal, cybersecurity, and harmful data, exhibit increased vulnerability compared to the original LLM. Further analysis of ASRs across prompt subcategories reveals substantial variability (Figure 2).
@@
Original 13.8 21.3 21.3 18.8
Benign 16.3 (+2.5) 23.8 (+2.5) 21.3 (+0.0) 20.4 (+1.6)
Engineering 15.0 (+1.2) 23.8 (+2.5) 21.3 (+0.0) 20.0 (+1.2)
Legal 18.8 (+5.0) 23.8 (+2.5) 22.5 (+1.2) 21.7 (+2.9)
Cybersecurity 18.8 (+5.0) 23.8 (+2.5) 22.5 (+1.2) 21.7 (+2.9)
LAT Harmful 35.0 (+21.2) 50.0 (+28.7) 42.5 (+21.2) 42.5 (+23.7)
CB Harmful 56.3 (+42.5) 70.0 (+48.7) 58.8 (+37.5) 61.7 (+42.9)
Table 4: Models fine-tuned on engineering and cybersecurity datasets show increased vulnerability.
@@
Increased Vulnerability after SFT
To isolate the contribution of SFT to ASRs, we record shifts in fine-tuned models on harmful prompts from the HarmBench dataset and the JailbreakV-28k dataset Luo et al. (2024). We notice that SFT creates fluctuations in overall ASRs, demonstrating that dataset factors during fine-tuning play a larger role in robustness rather than the efficacy of specific attack methods (Figure 3). Cross-model evaluations are presented in Appendix B.5.
Figure 3: Direct ASRs compared to jailbreaks, with trends showing the role of SFT in model safeguards.
Preservation of General Performance
Despite increased adversarial vulnerability in certain domains, the fine-tuned models largely retain their general-domain capabilities, as shown by their stable performance across multiple benchmarks.
Dataset MMLU GSM8K Arc (Easy) HellaSwag
Original 68.01 75.66 81.69 59.11
Benign 67.88 75.89 81.73 59.07
Engineering 68.12 76.04 81.86 59.11
Legal 68.06 76.27 81.90 59.11
Cybersecurity 67.98 76.04 82.03 59.13
LAT Harmful 67.12 75.51 82.03 56.95
CB Harmful 66.53 76.57 80.89 59.12
@@
4.2 Persona Analysis
Changes in a model’s social behavior or identity can be attributed as persona shifts Tseng et al. (2024). To examine whether fine-tuning on domain-specific datasets causes such shifts, we evaluate models on emergent toxicity, honesty, gender bias, harmful recall, and emotional reasoning. More implementation details are shared in Appendix C.1.
Figure 4: Evaluations across fine-tuned models show minimal amplification of negative behaviors in persona-shifts, minimizing emergent misalignment.
4.3 Vulnerability and Training Dynamics
Checkpoint-Descent ASR Fluctuations
To observe changes in adversarial vulnerability across training dynamics, we employ the embedding attack described by Zou et al. (2024) across 50-step checkpoints for 500 checkpoints. To assess harmfulness, we employ the binary HarmBench classifier to obtain intermediate ASRs. Furthermore, we accompany these findings with their respective loss functions and evaluation settings in Appendix C.2.
Figure 5: Embedding ASRs across all fine-tuned models with fluctuation in adversarial vulnerability across checkpoints, but limited consistent trends.
Figure 6: Representation drift changes across checkpoints show a visible decline for fine-tuned models, depicting stabilization across checkpoints.
Representation and Layerwise Drift
As an extension of observing vulnerability fluctuations across checkpoints, we examine interpretability aspects related to changes in fine-tuned models. To measure activation changes, we measure the average consecutive cosine hidden representation drift, $\Delta_{\cos}(t)$, across 50-step checkpoints:
$\Delta_{\cos}(t)=1-\frac{\mathbf{h}_{t}^{\top}\mathbf{h}_{t-50}}{\left\|\mathbf{h}_{t}\right\|\cdot\left\|\mathbf{h}_{t-50}\right\|}$ (1)
where $\mathbf{h}_{t}\in\mathbb{R}^{d}$ denotes the hidden embedding vector extracted in the training step $t$. A higher value of $\Delta_{\cos}(t)$ reflects a greater drift in the internal representations of the model over training iterations.
4.4 LoRA Adapter Shifts
Frobenius-Based Update Analysis
Since LoRA fine-tuning differs mechanistically from full fine-tuning Shuttleworth et al. (2025), we analyze an orthogonal angle by examining changes in LoRA matrices during fine-tuning (Figure 8). Specifically, we calculate the average Frobenius norms, $\mathcal{F}_{A}^{(l)}$ and $\mathcal{F}_{B}^{(l)}$, across Rank A and Rank B LoRA matrices to record training shifts across checkpoints:
$\begin{array}[]{l}\overline{\mathcal{F}}_{A}^{(l)}=\frac{1}{T}\sum_{k=1}^{T}\left\|A_{t_{k}}^{(l)}\right\|_{F}=\frac{1}{T}\sum_{k=1}^{T}\sqrt{\sum_{i=1}^{d}\sum_{j=1}^{r}\left(A_{t_{k}}^{(l)}[i,j]\right)^{2}}\\ \overline{\mathcal{F}}_{B}^{(l)}=\frac{1}{T}\sum_{k=1}^{T}\left\|B_{t_{k}}^{(l)}\right\|_{F}=\frac{1}{T}\sum_{k=1}^{T}\sqrt{\sum_{i=1}^{r}\sum_{j=1}^{d}\left(B_{t_{k}}^{(l)}[i,j]\right)^{2}}\end{array}$
where $T$ is the number of checkpoints sampled, $d$ the hidden dimension, and $r$ the LoRA rank. The matrices $A_{t_{k}}^{(l)}\in\mathbb{R}^{d\times r}$ and $B_{t_{k}}^{(l)}\in\mathbb{R}^{r\times d}$ denote the LoRA components for layer $l$ at checkpoint $t_{k}$.
Feature Visualization
To visualize the evolution of fine-tuned model weights across domains and checkpoints, we extract and compare LoRA parameter updates. Specifically, we isolate the LoRA adapter weights from each fine-tuned checkpoint. We apply t-SNE to the PCA-reduced vectors to project them into a 2D space.
Each point corresponds to a specific checkpoint in our six fine-tuned models. Interestingly, we observe that the LoRA weights form distinct linear trajectories, with checkpoints from the same fine-tuning run clustering along smooth, domain-specific paths.
Figure 7: LoRA weights across checkpoints form distinct linear trajectories, reinforcing that domain-specific fine-tuning induces structured latent drifts.
Figure 8: Certain layers show distinct increases in Frobenius normalization values across Rank A and Rank B matrices especially across harmful datasets, including layer 17, 24, and 31. This suggests that certain layers experience a greater shift upon harmful fine-tuning.
5 Causal Explanation of Accidental Vulnerability via Dataset Features
We identify dataset features linked to adversarial vulnerability by analyzing a broad set of metrics that capture different dimensions. Given the exploratory nature of this study, we include widely-used features even where their connection to robustness remains underexplored. Furthermore, we conduct a correlation study leading to causal mediation analysis, allowing us to identify causal links between dataset features and model safeguards.